OpenStudy (anonymous):

9^12/9^3 A. 1^4 B. 9^15 C. 9^4 D. 9^9

OpenStudy (anonymous):

\[\frac{ 9^{12} }{ 9^3 } = 9^{12-3} = 9^9\] The answer is D.